OpenStudy (mokeira):

I am sure my answer is correct although I was told it is wrong...anyone please confirm. (4+3i)*(7-i) My answer= -3i^2+17i+28

OpenStudy (anonymous):

Just a minor mistake.. :(

OpenStudy (agreene):

remember what \(i^2\) is

OpenStudy (mokeira):

oooh...so i am wrong :(

OpenStudy (anonymous):

Sorry, not mistake but you have forgot to do further.. :)

OpenStudy (mokeira):

\[i ^{2}= i \times i\]

OpenStudy (anonymous):

We know that \(i^2 = i \cdot i\) but what is the value of \(i^2\) ??

OpenStudy (mokeira):

i dont know..please explain

OpenStudy (anonymous):

Oh God what I am doing today.. :) -1 there..

OpenStudy (mokeira):

oook...so i= -1?

OpenStudy (anonymous):

\[\large \color{green}{i^2 = -1}\]

OpenStudy (agreene):

since i is an imaginary number that is defined by \(i^2=-1\) look at your answer and you will see what else you need to do.

OpenStudy (anonymous):

\(i = \sqrt{-1}\).. :)

OpenStudy (mokeira):

is it like a rule ?

OpenStudy (agreene):

yes

OpenStudy (mokeira):

ooooooh....i never knew that one

OpenStudy (anonymous):

As long as you are studying complex numbers, then it must be a rule.. :)

OpenStudy (mokeira):

Now i see why my answer was half way done!

OpenStudy (agreene):

as @waterineyes pointed out: \[i=\sqrt{-1}\] thus \[i^2=\sqrt{-1}^2 = -1\]

OpenStudy (anonymous):

@Mokeira are you studying Complex Numbers as your topic??

OpenStudy (mokeira):

@waterineyes I am!....teaching myself though that is why I was stuck

OpenStudy (anonymous):

I mean to say that what topic you are doing?

OpenStudy (mokeira):

Operations on Complex Numbers

OpenStudy (mokeira):

Sub topic in calculus I think

OpenStudy (anonymous):

Then you must know this : Real numbers cannot define some numbers like : Negative Numbers under Square roots.. So, we have made another number set called Complex Numbers set. Complex Number itself starts from : \[i = \sqrt{-1}\] As sqrt{-1} cannot be defined by real numbers so, we have chosen our set as Complex and decided to take -ve numbers under sqrt in terms of iota ie i..

OpenStudy (mokeira):

oooooooh wow! Thanks!!!!!! you are awesome. I have learnt something new

OpenStudy (anonymous):

Like : \[\large \color{green}{\sqrt{-16} = \sqrt{16} \cdot \sqrt{-1} = 4i}\]

OpenStudy (mokeira):

ok wait.... why not -4i?

OpenStudy (anonymous):

This way you can, any number, write in terms of i..

OpenStudy (agreene):

on the converse, squaring complex parts are done as such: \[bi^2=-b^2\] where b is a real number

OpenStudy (anonymous):

Why?? \(\sqrt{-1}\) is \(i\).

OpenStudy (mokeira):

now I get it....!

OpenStudy (mokeira):

@waterineyes @agreene You guys ROCK!!! thank youuuuu I have learnt soo much

OpenStudy (anonymous):

Keep on learning.. :)

OpenStudy (agreene):

np, that's why im here lol

OpenStudy (agreene):

so, getting back to your problem: \[-3i^2+17i+28\] \[3^2+17i+28\] and so on

OpenStudy (anonymous):

Only \(3\) there..

OpenStudy (anonymous):

Just a typing mistake nothing else.. :)

OpenStudy (mokeira):

so can you guys also explain to me how The equation \[ax ^{2} + bx + c = 0\] can be rewritten (when a is not 0, after dividing by a) as \[(x+\frac{ b }{ 2a }^{2})=\frac{ b ^{2}-4ac }{ 4a ^{2} }\]

OpenStudy (anonymous):

Yes it can be written as like you said.. :)

OpenStudy (mokeira):

i dont understand how you change from the equation to the fraction?

OpenStudy (anonymous):

To scare you.. :P

OpenStudy (mokeira):

lol... so its also like another rule?

OpenStudy (anonymous):

No, it can be achieved by some mathematical steps.. :)

OpenStudy (mokeira):

how...i have tried and cant figure it out

OpenStudy (anonymous):

Have you heard of "Completion of Square Method" ??

OpenStudy (mokeira):

it says divide by a

OpenStudy (anonymous):

So what loss you are bearing by dividing by a?? Just do it friend.. :P

OpenStudy (mokeira):

No I havnt...let me read about it then I try again

OpenStudy (anonymous):

Okay jokes apart... :) let us do it now :)

OpenStudy (mokeira):

Actually I know the "Completion of square method"

OpenStudy (anonymous):

Can you divide by a first?? As you are saying.. :)

OpenStudy (mokeira):

ok..let me divide by a

OpenStudy (anonymous):

If you know then it will solve your problem.. :)

OpenStudy (anonymous):

Oh my God @UnkleRhaukus uncle rocks is here only.. :) How are you Uncle Rocks??

OpenStudy (unklerhaukus):

hi !

OpenStudy (mokeira):

\[\frac{ a ^{2+bx+c} }{ a }=a+\frac{ bx }{ a }+\frac{ c }{ a }\] Am I right?

OpenStudy (anonymous):

Hi, Really I am very happy to see you here.. :) Long time..

OpenStudy (anonymous):

You should check it once again.. :)

OpenStudy (anonymous):

You have to divide Left Hand Side as well as Right Hand side by a..

OpenStudy (mokeira):

the equation to be divided is \[a ^{2}+bx+c\] there is no right and left hand side

OpenStudy (mokeira):

sorry..not an equation

OpenStudy (anonymous):

If no right hand side, then why are you saying it as equation??

OpenStudy (mokeira):

not an equation..i corrected myself..

OpenStudy (anonymous):

Okay, I am late.. :P

OpenStudy (mokeira):

oh wait!...its =0

OpenStudy (anonymous):

You are having an equation, Okay?/ \(ax^2 + bx + c = 0\)

OpenStudy (anonymous):

Are you alright @Mokeira ?? Something wrong??

OpenStudy (mokeira):

so even after you divide by a it will still be =0

OpenStudy (anonymous):

Yes, but check for Left hand side once again.. :)

OpenStudy (anonymous):

Can you clearly write down what have you got after dividing by a, write if fully with LHS and RHS like an equation.. :)

OpenStudy (mokeira):

ok...sure. Dont get tired lol :)

OpenStudy (anonymous):

He he he he.. Who is getting tired?? :)

OpenStudy (mokeira):

\[a ^{2}+bx+c=0\] \[\frac{ a ^{2}+bx+c=0 }{ a }\] \[a+\frac{ bx }{ a }+\frac{ c }{ a }=0\] Is this right

OpenStudy (anonymous):

I hope you are not kidding this time.. Are you hungry??

OpenStudy (anonymous):

Where is \(x^2\) ??

OpenStudy (mokeira):

we are dividing by a right? How and where does x^2 come from?

OpenStudy (anonymous):

\[\frac{ax^2 + bx + c}{a} = 0 \implies x^2 + \frac{b}{a}x + \frac{c}{a} = 0\]

OpenStudy (anonymous):

Dear, you should check your equation first.. :) See it with your eyeballs for 5 minutes.. :P

OpenStudy (mokeira):

lol! I always omit!!!! i guess im always in a hurry. my bad! I get it now! Clumsy me

OpenStudy (mokeira):

lol...thanks! thank u for your time too

OpenStudy (anonymous):

can you go ahead with this now? Just use Completing the square method you will reach there where you want.. :)

OpenStudy (anonymous):

It is okay.. :) You are welcome dear.. :)

OpenStudy (anonymous):

I have to go, if you have any doubt, then @UnkleRhaukus will help you here from nowonwards.. Keep doing more and more.. :)

OpenStudy (mokeira):

ok! thanks bye

OpenStudy (anonymous):

See you later @UnkleRhaukus :)

OpenStudy (unklerhaukus):

see you next time @waterineyes !